In a scale drawing of a playground, the total lenght is 5 cm. The actual length is 45 feet. If the actual width of the playground is 60 feet what is its width on the drawing?
Answer by rfer(16322)

5/45=x/60
45x=300
x=300/45
x=6.7 cm